HETAV PATEL

Grade 11,

A gas is enclosed in a vessel at a pressure of 2.5 atm. Due to a leak in the vessel, after some time the pressure is reduced to 2 atm, temperature remains unchanged. The percentage of gas that has leaked out is

 A gas is enclosed in a vessel at a pressure of 2.5 atm. Due to a leak in the vessel, after some time the pressure is reduced to 2 atm, temperature remains unchanged. The percentage of gas that has leaked out is

Grade:12th Pass

1 Answers

Mujahid Ahmed
12 years ago

∝ n

let no. of moles in first case be n. Assuming x moles leak out,

25/20 = n//n-x

25x = 5n

x/n = 1/5

(x/n).100 = 20%